OpenStudy (anonymous):

6x+6+5x−7=54 | Solve for x Answer

OpenStudy (anonymous):

u need help?

OpenStudy (crashonce):

ok shift all xs to one side and numbers to other

OpenStudy (anonymous):

put x for one side and other numbers to another

OpenStudy (anonymous):

yes ido

OpenStudy (anonymous):

x= 5

OpenStudy (anonymous):

i just want to know how to get it x=5

OpenStudy (ashes.boom):

combine the like terms (6x and 5x) and then the two coefficients

OpenStudy (anonymous):

6x+5x = 54+7 - 6

OpenStudy (josedavid):

WHAT GRADE ARE U IN

OpenStudy (anonymous):

11x = 55 x = 55 / 11

OpenStudy (josedavid):

JK DONT TAKE THAT THE WRONG WAY

OpenStudy (anonymous):

and why its a wrong way?

OpenStudy (josedavid):

JUST ASKING

OpenStudy (anonymous):

this was a geometry equation btw

OpenStudy (anonymous):

10th

OpenStudy (josedavid):

OKAY